OpenStudy (anonymous):

Highest and least value of 4x^2 in the the inequation

OpenStudy (anonymous):

sry wait typing the question

Parth (parthkohli):

I've seen all of these questions in my coaching module. o.O

Parth (parthkohli):

Since No.Name isn't here, I'll type the question out...

Parth (parthkohli):

\[\dfrac{6x^2 - 5x - 3}{x^2 - 2x + 6} \le 4\]

Parth (parthkohli):

What did you get for the solution set of \(x\)?

OpenStudy (anonymous):

My cputer hanged , sry yes that is the question

Parth (parthkohli):

OK, your first step is to solve the inequality...

OpenStudy (anonymous):

I am stuck on this step \[\huge \frac{ 2x^2 +3x-27 }{ x^2-2x+6 } <=0\]

OpenStudy (anonymous):

They can't be perfectly factorized , o have I done a very silly mistake

Parth (parthkohli):

The denominator is always positive, so you can just remove it from the inequality.

OpenStudy (anonymous):

yeah \[\huge2x^2+3x-27 \le 0\]

Parth (parthkohli):

Great, factorize.

OpenStudy (anonymous):

one oot is in decimal;s

Parth (parthkohli):

Yeah.\[2x^2 + 9x - 6x - 27 = x(2x + 9) - 3(2x + 9) = (x-3)(2x+9) \le 0\]\[x \in \left(-\frac{9}{2},3\right)\]

Parth (parthkohli):

Now is the fun part.

OpenStudy (anonymous):

Actually \[\huge [-4.5 , 3]\], they are included

Parth (parthkohli):

Argh, OS is buggy. Yeah - I was correcting that.

OpenStudy (anonymous):

but weird answer is given

Parth (parthkohli):

No...

Parth (parthkohli):

Do you know how \(|x|^2 = x^2\)?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

I am getting 36 and 81

Parth (parthkohli):

So if a number has the highest modulus, then its square will be the highest. Which means that -4.5 has a bigger modulus than 3, so it has the highest square in the solution set of x, meaning that it has the highest four times the square...

Parth (parthkohli):

Similarly the number with the least modulus in the solution set of x is 0, so it is supposed to have the least square, and also the least four-times-the-square.

Parth (parthkohli):

You got the 81 right.

OpenStudy (anonymous):

oh fu**k

Parth (parthkohli):

You mean fu*k.

OpenStudy (anonymous):

wtf man , I didn't realize that , I plugged in the exteme values lol

Parth (parthkohli):

lol

OpenStudy (anonymous):

0 is a cruel number thank you , !

Parth (parthkohli):

Ancient Indians are coming to kill you. :P

OpenStudy (anonymous):

rofl

OpenStudy (ikram002p):

0 :D is my favorite nothing :P

OpenStudy (anonymous):

because you get it on your exam am I right?

OpenStudy (ikram002p):

xD lol

OpenStudy (anonymous):

Nothing is your favourite yet something's you favourite because Hamiltonian of thoughts can never be 0 @ikram002p

OpenStudy (anonymous):

just jokin, hehe

OpenStudy (ikram002p):

ik udit :P but still a nice joke